City A and City B are 440 miles apart. If Jack leaves city A traveling 50mph and Jill leaves city B traveling 60mph, how long will they travel if they drive toward each other?
Answer by ankor@dixie-net.com(22740)

City A and City B are 440 miles apart. If Jack leaves city A traveling 50mph and Jill leaves city B traveling 60mph, how long will they travel if they drive toward each other?
:
Let t = traveling time when they meet
:
We know that when they meet, their combined travel distance will be 440 mi
:
Dist = speed * time
:
jack's dist + Jill's dist = 440
50t + 60t = 440
110t = 440
t = 440/110
t = 4 hrs
:
:
Check: 4(50) + 4(60) = 440